OpenStudy (anonymous):

1/5 of students in a school have no brothers or sisters. Of the remaining students, 40% have 1 brother or sister. What percent of all students have more than 1 brother or sister?

OpenStudy (anonymous):

hi tutor

OpenStudy (anonymous):

one fifth is 20%

OpenStudy (anonymous):

via \[\frac{1}{5}=.2=20\%\]

OpenStudy (anonymous):

so 20% have no siblings, 40% have one leaving....

OpenStudy (anonymous):

so 40% is the result?

OpenStudy (anonymous):

is it equation : 20%+40%+X= 100% THEN X = 40 IS THE RESULT?

OpenStudy (anonymous):

u still there?

OpenStudy (anonymous):

no it's 40% of the remaining, not of the total. 1-.8*.4=.68

OpenStudy (anonymous):

yes what is left is 40%

OpenStudy (anonymous):

???

OpenStudy (anonymous):

you have accounted for 60% of the total leaving 40% unaccounted for. that is the answer.

OpenStudy (anonymous):

100 students 1/5 don't have any which means only 80 have at least one of those 80, 40% (which is 32) have only one 80-32=68

OpenStudy (anonymous):

80-32=48 sorry :-)